Consider a bet where you have 50% chance of winning $40, a 30% chance of winning $60, and a 20% chance of winning $150
a. What is the expected payoff of this bet?
b. What is the value of the bet to someone with log utility and an initial wealth of $100? c. Is the value of the same bet any different to someone who also has log utility but an initial wealth of $200?
Answer #1

A) expected payoff EC = .5*40 + .3*60 + .2*150

= 20+18+30 = 68

B) U = log(w)

Then value of the bet is the maximum willingness to pay to avoid any riskiness in payoffs  

= Initial wealth - certainty equivalent

So EU = .5*log(40)+.3*log(60)+.2*log(150)

= 1.78

So for CE :

log(CE)= EU = 1.78

Thus CE = 60.256

So value = 100-60.256

= $ 39.744

c) yes value will be different

= 200-60.256

= $ 139.744